OpenStudy (anonymous):

How many total oxygen (O) atoms are present in the reactants side of the equation K2O + H2O → 2 KOH?

OpenStudy (australopithecus):

\[K_2O\] Has 1 oxygen \[H_2O\] Has 1 Oxygen 1 + 1 = 2 Therefore the reactant side has 2 oxygen atoms Also the equation is balanced because there are 2 oxygen atoms on the product side as you get 2 times KOH, thus you get 2K, 2O, 2H, which you have on the reactant side as well

OpenStudy (lgbasallote):

correct me if im wrong, but i believe the 2 you're referring to is the number of moles of Oxygen. It is true that oxygen has 2 moles in the reactant side. However, as for the atoms, you need to multiply it by avogrado's constant.. \[\Large \implies 2 \; \text{mole}\times \frac{6.022 \times 10^{23}\; \text{atoms}}{\text{moles}}\] \[\Large \implies 2 \; \cancel{\text{moles}}\times \frac{6.022 \times 10^{23}\; \text{atoms}}{\cancel{\text{moles}}}\] does that help?

OpenStudy (anonymous):

No